The width of a rectangle is 6 centimeters less than the length.The perimeter is
64 centimeters.Find the width.
One angle of a triangle is 20 degrees.The second angle's measurement is four times the third.What is the measure of the third angle?
The sum of two angles in a triangle totals 62.5 degrees.What is the other angle's measure?

Answer by rfer(16322)

A)
2x+2(x-6)=64
4x-12=64
4x=76
x=19
x-6=13
B)
20+4x+x=180
5x=160
x=32 degrees
4x=128 degrees
C)
180-62.5=117.5 degrees